If is estimated that 30% of all students who fall under the Business Economics and Management department at Richfield have drivers licence. What is the probability of this in a sample of 10 students


Exactly 4 students have drivers license


1
Expert's answer
2021-05-11T03:55:08-0400

"n=10, p=0.3,q=1-0.3=0.7"


Probability Exactly 4 students have driver license


"=^{10}C_4p^4q^{10-4}\\

 =^{10}C_4(0.3)^4(0.7)^6=0.20"